Andrea has $4.15 in dimes and quarters. She has 3 more dimes than quarters. How many of each coin does she have?
Answer by nerdybill(7384)

Andrea has $4.15 in dimes and quarters. She has 3 more dimes than quarters. How many of each coin does she have?
.
Let q = number of quarters
then
q+3 = number of dimes
.
.25q + .10(q+3) = 4.15
.25q + .10q+.30 = 4.15
.35q + .30 = 4.15
.35q = 3.85
q = 11
.
dimes:
q+3 = 11+3 = 14
.
Answer: 11 quarters and 14 dimes